### v1

# RIS-LAD: A Benchmark and Model for Referring Low-Altitude Drone Image Segmentation

Abstract
Referring Image Segmentation (RIS), which aims to segment specific objects based on natural language descriptions, plays an essential role in vision-language understanding. Despite its progress in remote sensing applications, RIS in Low-Altitude Drone (LAD) scenarios remains underexplored. Existing datasets and methods are typically designed for high-altitude and static-view imagery. They struggle to handle the unique characteristics of LAD views, such as diverse viewpoints and high object density. To fill this gap, we present RIS-LAD, the first fine-grained RIS benchmark tailored for LAD scenarios. This dataset comprises 13,871 carefully annotated image-text-mask triplets collected from realistic drone footage, with a focus on small, cluttered, and multi-viewpoint scenes. It highlights new challenges absent in previous benchmarks, such as category drift caused by tiny objects and object drift under crowded same-class objects. To tackle these issues, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN). Rather than uniformly injecting all linguistic features, SAARN decomposes and routes semantic information to different stages of the network. Specifically, the Category-Dominated Linguistic Enhancement (CDLE) aligns visual features with object categories during early encoding, while the Adaptive Reasoning Fusion Module (ARFM) dynamically selects semantic cues across scales to improve reasoning in complex scenes. The experimental evaluation reveals that RIS-LAD presents substantial challenges to state-of-the-art RIS algorithms, and also demonstrates the effectiveness of our proposed model in addressing these challenges. The dataset and code will be publicly released soon at: https://github.com/AHideoKuzeA/RIS-LAD/.
Index Terms:
referring image segmentation, unmanned aerial vehicle, drone, aerial image analysis, benchmark dataset.
I Introduction
Low-Altitude Drones (LAD), which typically operate below 200 meters, have become widely utilized in real-world applications due to their flexible deployment and high versatility [1, 2, 3]. This trend has sparked increasing research interest in vision tasks under LAD scenarios, leading to the development of benchmarks such as VisDrone [4] and UAV123 [5], which support object detection, tracking and other related tasks [6, 7, 8]. Recently, visual understanding in LAD scenarios [9, 10] has attracted increasing interest, especially in multi-modal tasks. Among them, Referring Image Segmentation (RIS) [11, 12, 13] is a fundamental perception task that aims to segment objects from images based on descriptions. Incorporating RIS into LAD systems enables LADs to better support a wider range of practical applications.
Figure 1: Illustration of the challenges in Referring Low-Altitude Drone Image Segmentation (RLADIS). The method shown in (b) is RefSegRS [14], one of the current SOTA approaches for RRSIS.
However, existing RIS research primarily focuses on conventional scenes or high-altitude remote sensing imagery and lacks systematic investigations tailored to LAD imagery. Although previous Referring Remote Sensing Image Segmentation (RRSIS) studies [15, 14, 16] have provided valuable insights, they are mostly based on imagery taken from satellites or helicopters. As illustrated in Fig. 1(a), Referring Low-Altitude Drone Image Segmentation (RLADIS) presents several unique challenges compared to RRSIS: (1) diverse viewpoints resulting from lower and more flexible camera positions; (2) diverse illumination conditions; and (3) smaller and more densely distributed objects. These challenges introduce a substantial domain gap, which makes it difficult for existing RRSIS methods to generalize effectively to RLADIS tasks.
To bridge this gap, we propose RIS-LAD, the first dataset designed for fine-grained RLADIS. In addition, a semi-automatic annotation pipeline is also developed to ensure high-quality and efficient mask generation. The final dataset consists of 13871 image-text-mask triples and serves as a benchmark for evaluating RLADIS approaches. Furthermore, our analysis reveals two key challenges that hinder the direct adaptation of RRSIS methods to RLADIS: category drift and object drift. As illustrated in Fig. 1(b), category drift arises when the object occupies only a small region of the image, leading the model to mistakenly focus on larger, semantically similar objects. Object drift is caused by a high density of instances in the same category, which makes it difficult for the model to correctly identify the object described by the expression.
Motivated by the issues of category and object drift in RLADIS, we introduce the Semantic-Aware Adaptive Reasoning Network (SAARN), which decouples linguistic features and injects them at appropriate stages of the network to ensure semantic alignment. In contrast to existing decoupling-based approaches that inject all linguistic components uniformly across modules [17, 18], SAARN follows a more targeted strategy. Specifically, the Category-Dominated Linguistic Enhancement (CDLE) module injects only class-level linguistic features in the encoder stage. This enables early visual representations to be aligned with accurate category semantics. Meanwhile, the global linguistic features are selectively integrated via class-guided gating, which further reinforces semantic consistency. In the multi-scale fusion stage, the Adaptive Reasoning Fusion Module (ARFM) performs scale-aware enhancement by dynamically weighting semantic cues across multi-scale features. This mechanism simulates a reasoning process that guides the model to infer the most semantically consistent instance among densely packed objects of the same category.
The main contributions of this paper are as follows:
  • We introduce RIS-LAD, the first Referring Image Segmentation (RIS) dataset tailored for Low-Altitude Drone (LAD) scenarios, addressing the lack of fine-grained annotated data in this domain. It features diverse LAD scenes with rich textual descriptions and high-quality segmentation masks.
  • To address category drift and object drift in RLADIS, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN), which emphasizes semantic category alignment and selectively integrates the most relevant linguistic features at appropriate feature scales.
  • Through comprehensive experiments, we investigate the challenges in RLADIS and construct a benchmark tailored for this task. Meanwhile, the proposed SAARN achieves state-of-the-art performance on the RIS-LAD benchmark.
II RELATED WORKS
II-1 Referring Image Segmentation (RIS)
Referring Image Segmentation (RIS) aims to achieve pixel-level localization of objects in images based on natural language expressions [19]. To address key limitations in RIS, recent research has explored advanced strategies. IteRPrimE [12] refines visual activation maps through an iterative Grad-CAM refinement strategy and adopts a primary-word emphasis module to improve the identification of key semantic components in language expressions. VATEX [20] leverages a CLIP-based prior module to generate heatmaps, which enforces semantic consistency and improve contextual alignment between language and vision. ASDA [21] utilizes adaptive selection and dual-alignment methods to enable adaptive alignment of visual and linguistic features. DeRIS [22] decomposes the RIS task into perception and cognition components, and employs a loop-back synergy mechanism to achieve bidirectional enhancement between the two modules. However, while these approaches perform well on conventional images, their effectiveness significantly declines when applied to the LAD imagery, due to substantial domain gaps between the two types of imagery.
II-2 Referring Remote Sensing Image Segmentation (RRSIS)
RRSIS is a domain-specific extension of the RIS task for remote sensing images, first introduced by [23]. Existing RRSIS benchmarks [24, 25, 26] are mainly constructed from high-altitude sources such as Google Earth and GF-2. Recent methods [27, 28] have advanced the performance of RRSIS through different strategies. FIANet [17] disentangles various components of the referring expression and incorporates a text-aware multi-scale enhancement module to improve multi-modal discrimination. RMSIN [25] designs rotation-aware modules to precisely segment objects based on the unique properties of remote sensing imagery. RSRefSeg integrates a CLIP-based [29] module to capture implicit visual activations and uses them as guidance prompts for SAM [30]. DiffRIS [31] leverages the semantic understanding capabilities of pretrained text-to-image diffusion models to improve cross-modal alignment. Although some benchmarks have considered low-altitude scenarios (Li and Zhao 2025), drone imagery still presents unique challenges. These include tiny object sizes and high object density, which result from low flying altitudes and diverse viewpoints, significantly complicating both semantic grounding and visual recognition.
Figure 2: Illustration of the semi-automatic annotation pipeline for RIS-LAD. Mask generation leverages SAM-2 to extract instance-level segmentation masks. Referring expression generation employs Qwen2.5-VL to produce initial descriptions conditioned on location markers and cropped object instances, with subsequent human correction.
III RIS-LAD Dataset
This section presents the construction pipeline of the proposed RIS-LAD dataset in detail.
III-1 Image Selection
RIS-suitable images are selected from the publicly available CODrone dataset [32], which is designed for LAD-oriented object detection. The selected images cover diverse real-world urban scenes under different lighting conditions and from various viewpoints. Following common procedures in constructing RRSIS datasets [26, 24, 33], the selected high-resolution images are divided into uniform $\times 1080$ 1080 tiles, thereby preserving fine-grained details. A manual screening step is conducted to ensure that the retained tiles exhibit representative characteristics of LAD imagery.
III-2 Object Selection
[…]
III-3 Mask and Expression Generation Pipeline
The Segment Anything Model 2 (SAM-2) [34] is utilized to generate high-quality instance masks in an automated manner. Specifically, oriented bounding boxes provided by the CODrone serve as precise prompts to guide the segmentation process.
As illustrated in Fig. 2, referring expressions are generated using the multi-modal large language model (MM-LLM) Qwen2.5-VL [35]. In order to capture both instance-level appearance and spatial attributes, each object is cropped and upsampled to improve visual clarity. The cropped instance and the corresponding full image, with a red bounding box serving as the location marker, are then fed into Qwen2.5. This dual-input strategy helps mitigate the challenges posed by small object sizes. These objects often hinder the MM-LLM’s ability to capture fine-grained visual features and generate accurate spatial descriptions when relying solely on raw coordinates. As shown in Fig. 3, we further visualize linguistic diversity through word clouds generated from the refined referring expressions. The complete prompting strategy is provided in the supplementary materials.
Figure 3: Word cloud for referring expressions in the RIS-LAD dataset
III-4 Manual Refinement and Dataset Splitting
All automatically generated annotations are carefully verified and refined to ensure strong semantic alignment among images, referring expressions, and masks. This process yields a total of 13,871 high-quality image-text-mask triplets. These samples are then split into training, validation, and test sets with a 7:1:2 ratio.
Dataset Image Source Shooting Angle Nighttime Scene RefDIOR [24] Google Earth fixed NWPU-Refer [33] Google Earth fixed RISBench [24] Google Earth, GF-2, JL-1 fixed RefSegRS [26] Helicopter (above 1000m) fixed RRSIS-D [25] Google Earth fixed RIS-LAD Low-Altitude Drone $30^{\circ}$$60^{\circ}$
TABLE I: Comparison of RIS-LAD and existing referring remote sensing image segmentation datasets.
Figure 4: Illustration of the proposed Semantic-Aware Adaptive Reasoning Network (SAARN). The framework comprises two core components: Category-Dominated Linguistic Enhancement (CDLE, top) and Adaptive Reasoning Fusion Module (ARFM, bottom). CDLE injects class-level linguistic features at the encoder stage to mitigate early-stage semantic misalignments. ARFM adaptively integrates multi-scale and linguistic features to reason about the correct object in scenes with dense object distributions. $\omega$ and $\varphi$ denote linear and nonlinear mappings, respectively.
III-5 Comparison of RIS-LAD and RRSIS datasets.
As summarized in Table I, RIS-LAD exhibits distinct characteristics that distinguish it from existing RRSIS datasets. First, in contrast to RRSIS datasets that rely on fixed top-down viewpoints, RIS-LAD contains imagery captured from low-altitude drones at oblique angles between $30^{\circ}$ and $60^{\circ}$, thus offering greater variation in perspective and object appearance. Secondly, RIS-LAD uniquely includes nighttime scenes, enabling evaluation under challenging illumination conditions. Furthermore, it employs a semi-automatic annotation pipeline that integrates MM-LLMs with manual verification, thereby ensuring both annotation quality and efficiency.
IV Method
IV-A Overview
As illustrated in Fig. 4, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN), a framework designed to address category drift and object drift in Referring Low-Altitude Drone Image Segmentation (RLADIS). Built on RMSIN, SAARN incorporates two key components: the Category-Dominated Linguistic Enhancement (CDLE) and the Adaptive Reasoning Fusion Module (ARFM).
Given an input image, a Swin Transformer [36] encoder extracts visual features, while the referring expression is decomposed into global, class-level, and descriptive components, yielding three types of linguistic features: $l$, $c$, and $d$. For instance, in the expression ”The red car in the middle of the image parked in front of the white car,” the entire sentence corresponds to the global linguistic component; ”car” represents the class-level component; and the phrase ”red [masked] in the middle of the image parked in front of the white car,” excluding the category word, serves as the descriptive component.
At the encoder stage, CDLE injects class-level guidance through $c$ to align early visual representations with the correct category. Subsequently, it integrates $l$ to construct category-aware visual semantics. In the multi-scale fusion stage, ARFM employs an adaptive integration of pyramid fooled features with $l$, $c$, and $d$ across different scales. It assigns adaptive weights based on semantic alignment and spatial granularity, allowing each scale to emphasize the most relevant linguistic cues. The final fused representation is then passed to a mask decoder to generate the segmentation output.
IV-B Category-Dominated Linguistic Enhancement
In order to address category drift, we propose the Category-Dominated Linguistic Enhancement (CDLE) module, which introduces only class-level linguistic features to align early visual representations with accurate semantic categories. Descriptive linguistic components are deliberately excluded, as they may align with visually similar but incorrect objects, thereby worsening category drift. This module is integrated into each stage of the visual encoder to enable fine-grained cross-modal interaction.
The visual features, or the output from the previous encoder layer, are denoted as $x\in\mathbb{R}^{B\times HW\times D}$, where $B$ is the batch size, $HW$ is the flattened spatial resolution, and $D$ is the dimension of the feature. The class-level linguistic features $c\in\mathbb{R}^{B\times D_{b}\times N}$ are derived by extracting the category token from the referring expression and encoding it using a pre-trained BERT model [37]. $D_{b}$ and $N$ denote the hidden dimension and the number of linguistic tokens, respectively. Inspired by [38], we project both $x$ and $c$ into a shared embedding space to compute the scaled dot-product attention to align category cues and visual features:
$\alpha^{c}=\text{softmax}\left(\frac{\omega^{i}_{q}(x)\cdot\left(\omega^{c}_{k}(c)\right)^{\top}}{\sqrt{d}}\right)\cdot\omega^{c}_{v}(c),$ (1)
where each $\omega$ denotes a convolutional layer applied along the channel dimension, and $d$ is the scaling factor.
Although the category-guided attention $\alpha^{c}$ effectively captures semantic alignment between the input and class-level linguistic cues, it tends to over-emphasize category semantics while suppressing other informative visual contexts. To offset this, a residual gate mechanism is employed, which facilitates adaptive weighting between $\alpha^{c}$ and the original feature $x$:
$z^{c}=\varphi^{ic}_{o}(\omega^{ic}_{w}(\alpha^{c})\odot\varphi_{m}^{i}(x)),$ (2)
$f^{c}=x+z^{c}\cdot\varphi_{f}^{ic}(z^{c}),$ (3)
where $\varphi_{m}^{i}(\cdot)$ and $\varphi^{ic}_{o}(\cdot)$ denote the convolutional layer followed by GELU. And $\varphi^{ic}_{f}(\cdot)$ denotes a two-layer linear projection with ReLU [39] and Tanh activations.
Given the global linguistic features $l\in\mathbb{R}^{B\times D_{l}\times N}$ and the category-guided feature $f^{c}$, we again compute the scaled dot-product attention and residual gate,
$a^{l}=\text{softmax}\left(\frac{\omega^{ic}_{q}(f^{c})\cdot\left(\omega^{l}_{k}(l)\right)^{\top}}{\sqrt{d}}\right)\cdot\omega^{l}_{v}(l),$ (4)
$z^{l}=\varphi_{o}^{icl}(\omega^{icl}_{w}(\alpha^{l})\odot\varphi_{m}^{ic}(f^{c})),$ (5)
$f^{l}=f^{c}+z^{l}\cdot\varphi_{f}^{icl}(z^{l}),$ (6)
This design ensures that $l$ is selectively activated in regions corresponding to $c$, thus providing a safeguard mechanism for the fusion between $l$ and $f^{c}$. The output $f^{l}$ retains the discriminative semantics of the referring expression while maintaining strong alignment with the correct category. This effectively mitigates attention drift toward visually or semantically similar but incorrect instances.
Figure 5: The architecture of the Adaptive Reasoning Fusion Block. Three types of linguistic features are enhanced via Multi-Head Attention (MHA).
V Adaptive Reasoning Fusion Module
The Adaptive Reasoning Fusion Module (ARFM) has been proposed to address the problem of object drift induced by densely packed same-category instances in LAD scenarios. Inspired by the human perception process, in which reasoning typically begins with coarse spatial localization and gradually refines toward specific object attributes, ARFM adopts a progressive fusion strategy. It establishes a dynamic weighting scheme between the fused multi-scale features $X^{f}=\{x^{f}_{1},x^{f}_{2},x^{f}_{3},x^{f}_{4}\}$ and the linguistic features $l$, $c$, and $d$. Thus, the model can capture scale-specific semantic cues more effectively, enabling semantic-aware attention across spatial resolutions. For instance, fine-grained details, such as color or shape, can be better captured at higher spatial resolutions, while coarse spatial descriptors such as “top-left corner,” are more salient at lower resolutions.
$X^{f}$ is first aligned using pyramid pooling, which is then downsampled to a shared resolution and channel dimension, resulting in $X^{fp}\in\mathbb{R}^{B\times C^{\prime}\times H^{\prime}\times W^{\prime}}$. These features are then fed into the Adaptive Reasoning Fusion Block (ARFB), which performs multi-branch, cross-modal reasoning guided by linguistic features. As illustrated in Fig. 5, ARFB contains three parallel linguistic branches that attend to $c$, $d$ and $l$, respectively. $X^{fp}$ is projected into a shared embedding space and interacts with each linguistic feature. Formally, for each linguistic feature $s\in\{l,d,c\}$, we compute:
$q_{s}$ $=W^{q}_{s}X^{fp}+b^{q}_{s},\quad k_{s}=W^{k}_{s}s+b_{s}^{k},$ (7)
$v^{s}$ $=W_{s}^{v}s+b_{s}^{v},\quad\ \ \ \ \ \alpha_{s}=\text{MHA}(q_{s},k_{s},v_{s})$ (8)
[…]
Method Publication P@0.5 P@0.6 P@0.7 P@0.8 P@0.9 oIoU mIoU Val Test Val Test Val Test Val Test Val Test Val Test Val Test LAVT[38] CVPR 2022 33.95 31.67 31.16 27.04 26.29 23.02 19.41 17.55 9.89 8.58 44.03 41.97 32.25 30.14 ASDA[21] MM 2024 35.74 33.19 29.44 26.85 23.14 20.16 13.54 12.78 4.01 3.94 38.70 37.53 35.46 33.33 VATEX[20] WACV 2025 19.27 16.07 14.76 11.87 9.74 8.47 6.02 4.67 2.08 1.63 24.83 24.27 20.32 18.53 LGCE[26] IEEE TGRS 2024 28.22 26.38 24.07 21.82 19.99 17.55 15.11 13.32 7.88 7.06 41.72 40.75 27.68 26.17 FIANet[17] IEEE TGRS 2024 42.91 40.21 37.61 35.22 32.59 29.93 26.43 23.71 14.83 13.21 45.24 43.39 39.61 37.44 RMSIN[25] CVPR 2024 45.85 43.36 40.97 38.11 35.24 32.36 28.01 25.08 16.33 13.75 50.17 48.82 42.08 39.60 RSRefSeg[14] IGARSS 2025 47.35 44.73 44.48 40.21 39.76 36.30 35.17 30.91 26.58 21.46 50.04 47.71 43.92 41.16 CADFormer[41] JSTARS 2025 45.34 42.20 38.90 36.99 33.60 31.92 26.86 23.60 15.40 12.63 47.37 46.47 41.36 39.32 SAARN(ours) 47.06 45.02 43.12 39.34 38.25 33.37 31.81 26.24 19.27 15.31 51.54 49.60 44.30 41.67
TABLE II: Comparison with state-of-the-art RIS and RRSIS methods on the proposed RIS-LAD dataset.
To adaptively modulate the contribution of each semantic branch, a Scale Reasoning Gate (SRG) module based on $X^{fp}$ is introduced:
$[w_{l},w_{d},w_{c}]=\text{Softmax}(\text{SRG}(X^{fp}))$ (9)
[…]
$\alpha^{f}=\text{Fuse}(w_{l}\alpha_{l},\ w_{d}\alpha_{d},\ w_{c}\alpha_{c})$ (10)
[…]
Finally, $\alpha^{f}$ is combined with $X^{fp}$ via residual addition and further enhanced through a feed-forward network [40], before being passed into the subsequent scale-aware gate fusion module [25]. ARFM adaptively incorporates different linguistic features into the main semantic path, placing varying emphasis on each feature to enable it to focus on the regions to which it is most sensitive. This improves context consistency and the quality of object boundary prediction.
VI Experiments
VI-A Implementation Details
We utilize the pre-trained Swin Transformer [36] as the visual backbone to extract multi-scale image features. For linguistic encoding, we employ the BERT-base model [37] to obtain contextual embeddings of referring expressions. The model is trained on the proposed RIS-LAD dataset, specifically curated to evaluate performance in realistic LAD environments. The model is trained for 50 epochs using the AdamW optimizer, with an initial learning rate of $3\times 10^{-5}$ and a weight decay of 0.01. A polynomial decay schedule is employed to progressively reduce the learning rate. All experiments are conducted on four NVIDIA RTX 3080 GPUs, using a batch size of 2 per GPU.
For metrics, we report Precision@X (P@0.5–0.9), Overall Intersection-over-Union (oIoU), and Mean Intersection-over-Union (mIoU). As in previous RRSIS studies, oIoU and mIoU serve as the primary metrics for evaluating segmentation performance.
VI-B Comparison with SOTA Methods
As shown in Table II, the proposed SAARN achieves the best performance across both primary metrics: oIoU and mIoU. Although it is based on RMSIN, our approach improves oIoU from 50.17 to 51.54 and mIoU from 42.08 to 44.30 on the validation set. It further achieves a significant mIoU improvement of 2.07 on the test set. Notably, SAARN achieves substantial gains under stricter localization thresholds, including +18.0% in P@0.9 and +13.6% in P@0.8. These results demonstrate the effectiveness of the proposed method in improving semantic grounding and object discrimination in complex LAD scenarios.
Conventional RIS methods, such as ASDA and VATEX, face significant performance degradation under the LAD setting, mainly due to the small size of objects and the complex background. In particular, VATEX shows relatively limited performance, partially because it relies on a frozen CLIP image encoder for visual enhancement. This leads to negative transfer because of the difference in domain between LAD imagery and the CLIP pre-training distribution. This observation underscores the fundamental domain gap between LAD-RIS and conventional RIS.
RSRefSeg directly integrates CLIP and SAM, achieving strong results on Precision@0.5–0.9. However, qualitative results in Fig. 6 reveal that RSRefSeg tends to predict overly large masks that cover both target objects and irrelevant objects. Consequently, the high precision scores may not accurately reflect the model’s true localization ability, potentially providing a misleading indication of performance in the presence of excessive mask coverage. Compared to FIANet, which also employs linguistic feature decoupling and multi-scale integration, SAARN achieves significantly better performance, particularly on oIoU and mIoU metrics. We attribute this to the high density of same-category object distributions in LAD imagery, which exacerbates ambiguity during scale-level reasoning.
CDLE ARFM oIoU mIoU Val Test Val Test 49.77 48.32 42.08 39.60 51.31 49.70 43.97 40.68 49.82 49.28 43.31 41.02 51.54 49.60 44.30 41.67
TABLE III: Ablation study of CDLE and ARFM.
l c d oIoU mIoU Val Test Val Test 49.82 49.28 43.31 41.02 50.64 49.44 43.52 41.43 49.90 49.30 43.47 41.17 51.06 49.52 43.98 41.61 51.54 49.60 44.30 41.67
TABLE IV: Ablation study of linguistic components.
Figure 6: Qualitative comparisons between SAARN and the previous SOTA RMSIN and RefSegRS. The left part illustrates the predictions on tiny-object examples, where category drift easily occurs. The right part shows predictions in high-density scenarios, which are prone to object drift.
VI-C Ablation Study
We conduct ablation studies on the two core modules of SAARN, CDLE and ARFM, to validate their effectiveness. As shown in Table III, adding CDLE significantly improves mIoU. By reinforcing category semantics early in the process, CDLE helps the model focus on the correct object categories, resulting in better instance-level segmentation precision. Moreover, integrating ARFM significantly improves oIoU, increasing it from 49.82 to 51.54 on the validation set. This improvement stems from ARFM’s ability to distinguish densely packed objects across scales, enabling more accurate localization and reducing false activations.
VI-C1 Linguistic components ablation
As shown in Table IV, we further conduct an analysis of the linguistic components used in the ARFM module. Specifically, we evaluate the contributions of the global, class-level and descriptive linguistic feature $l$, $c$ and $d$. Adding $l$ increases the oIoU of the validation set from 49.82 to 50.64, indicating that $l$ provides a general semantic bias that facilitates the alignment of basic characteristics across scales. Interestingly, when $d$ is further introduced with $l$, performance drops slightly. The decrease suggests that, without class-level guidance, the model becomes more susceptible to semantic interference from fine-grained descriptive cues, which may disrupt category-consistent representation learning, especially at coarser resolutions. In contrast, introducing $c$ significantly improves performance on both oIoU and mIoU, with the validation oIoU increasing by 0.42. This highlights the critical role of $c$ in semantic alignment, as it provides strong category-specific constraints. Finally, combining all three features achieves the best overall performance. These results confirm that $c$ provides essential semantic structure, while $d$ contributes complementary fine-grained cues, leading to more accurate reasoning process.
VI-D Visualization
We perform a qualitative comparison with two of the SOTA RRSIS methods to provide a clearer understanding of our model’s effectiveness. As shown in Fig. 6, in typical LAD scenarios featuring tiny and densely packed objects, our model effectively identifies the correct object category. It also accurately distinguishes the target from nearby same-category instances, even under complex backgrounds and varying drone viewpoints. By contrast, selected RRSIS methods suffer from noticeable drift, including misclassification and confusion among densely clustered same-category objects.
VII Conclusion
In this work, we introduced RIS-LAD, the first fine-grained dataset designed for RLADIS. Through detailed analysis, we identified two core challenges specific to this setting: category drift and object drift, both of which significantly degrade the performance of existing RES and RRSIS models. To address these issues, we propose a Semantic-Aware Adaptive Reasoning Network (SAARN). SAARN disentangles linguistic features and injects them at different stages of the network. The CDLE module strengthens early-stage category alignment to reduce semantic ambiguity, while the ARFM performs adaptive multi-scale reasoning to suppress object drift in densely distributed scenes. The experimental results demonstrate the effectiveness of our approach, showing improvements over the SOTA RIS and RRSIS methods. Our work provides a strong foundation for multi-modal understanding in complex drone-based tasks.